OpenStudy (anonymous):

Factors of 4x^2y + 24x^2 – 24y – 144? Btw, the exponent in the first term is only 2, not 2y.

OpenStudy (anonymous):

= 4(x^2y + 6x^2 - 6y - 36) = 4(y(x^2 - 6) + 6(x^2 - 6)) = 4(x^2-6)(y + 6)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw